When 100 mL of 0.05 M formic acid is titrated with 0.05 M NaOH, what is pH at equivlanece point?
Use pKa of formic acid = 3.75
Answer #1

when yoomi af 0.05m formie acid titrated with 0.05m Nooh, it form a butten solution. pH = pka + log I salt og Hсоон + Na он —At equivalenu point number of moles of Nach = numbers of moles of Heooh acid. So, I salt) A coona = [acid] Acook.Тиш Coona as pH = pka + log (sl) и CooNu Tacid) A COOH - PH = Рka + о , P H = PKA Aeu pKA = 3:15 :. РН = 3-15 р,